The p.d applied to a 1 kΩ resistance in order that a current of 100 μA may flow is


1
Expert's answer
2021-08-24T14:47:23-0400

The Ohm's law says

"I=V\/R"

Hence, the potential difference

"V=I*R\\=100*10^{-6}\:\rm A*1000\: \Omega=0.1\:\rm V"